# Using Large Language Models for Template Detection from Security Event Logs

### v1

@@ VI Conclusion and Future Work
In previous studies, a generic wildcard notation (e.g., delimited-⟨$\langle$*$\rangle$ ∗ ⟩) has been used for denoting a variable part in log messages that match the template, and the current paper has adopted the same approach. However, such variable parts can have a different nature. For example, in Figure 1 (Section I), delimited-⟨$\langle$*$\rangle$ ∗ ⟩ can match IP addresses, port and process numbers, user names, etc. Instead of using a generic delimited-⟨$\langle$*$\rangle$ ∗ ⟩ notation, more specific notations for different data types (e.g., $\langle*:number\rangle$ or $\langle*:ipaddress\rangle$) would make detected templates significantly more informative. Whereas traditional algorithms would need regular expression-based event log preprocessing to achieve that goal (e.g., see [12, 5]), LLMs have the potential of detecting such templates from the original event log data, and studying approaches of finding these templates is a valuable further research direction.
@@
Finally, as mentioned in Section I, there is a shortage of publicly available recent security event logs, which would facilitate the cyber security-centric event log analysis research. Although this paper has published several such event logs, creating additional realistic security event log data sets and releasing them into the public domain remains an important task for future work.
Acknowledgment
This work was supported by the EXAI project of the Estonian Ministry of Education and Research (project number TK213U10).

### v3

@@ 6 Conclusion and Future Work
In previous studies, a generic wildcard notation (e.g., delimited-⟨$\langle$*$\rangle$ ∗ ⟩) has been used for denoting a variable part in log messages that match the template, and the current paper has adopted the same approach. However, such variable parts can have a different nature. For example, in Figure 1 (Section 1), delimited-⟨$\langle$*$\rangle$ ∗ ⟩ can match IP addresses, port and process numbers, user names, etc. Instead of using a generic delimited-⟨$\langle$*$\rangle$ ∗ ⟩ notation, more specific notations for different data types (e.g., $\langle*:number\rangle$ or $\langle*:ipaddress\rangle$) would make detected templates significantly more informative. Whereas traditional algorithms would need regular expression-based event log preprocessing to achieve that goal (e.g., see logcluster ; drain ), LLMs have the potential of detecting such templates from the original event log data, and studying approaches of finding these templates is a valuable further research direction.
@@
One of the limitations of the current study is the template detection performance analysis for a limited number of local LLMs. Since new LLMs are constantly released, evaluating the template detection capabilities of a larger number of local LLMs is an open future research topic.
Finally, as mentioned in Section 1, there is a shortage of publicly available recent security event logs, which would facilitate the cyber security-centric event log analysis research. Although this paper has published several such event logs, creating additional realistic security event log data sets and releasing them into the public domain remains an important task for future work.
Acknowledgements.
This work was supported by the Estonian Centre of Excellence in Artificial Intelligence (EXAI), funded by the Estonian Ministry of Education and Research grant TK213.